OpenStudy (anonymous):

2A+B=3C How fast is C increasing if A is decreasing at a rate of 0.190 M/s and B is decreasing at a rate of -9.5X10^-2 M/s?

OpenStudy (aaronq):

look at the stoichiometric coefficients

OpenStudy (aaronq):

\(-\dfrac{1}{2}\dfrac{d[A]}{dt}=\dfrac{1}{3}\dfrac{d[C]}{dt}\)

OpenStudy (anonymous):

so C would be decreasing at 1/3 the rate of A?

OpenStudy (aaronq):

no, you'd need to isolate the rate of change of C to find it's value \(-\dfrac{1}{2}\dfrac{d[A]}{dt}=\dfrac{1}{3}\dfrac{d[C]}{dt}\rightarrow \dfrac{d[C]}{dt}=\dfrac{3}{2}\dfrac{d[A]}{dt}\)

OpenStudy (anonymous):

ok, so it ends up being 0.285 Thanks!

OpenStudy (aaronq):

yep! no problem dont forget units